Circular No. 9285

Central Bureau for Astronomical Telegrams

Postal Address: Hoffman Lab 209; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbatiau@eps.harvard.edu  ISSN 0081-0304  cbat@iau.org

               URL http://www.cbat.eps.harvard.edu/index.html

Prepared using the Tamkin Foundation Computer Network

V1657 SCORPII = NOVA SCORPII 2017 = PNV J16521887–3754189

S. Nakano, Sumoto, Japan, reports the discovery of an apparent nova (mag 11.7) by H. Nishimura (Kakegawa, Shizuoka-ken, Japan) on three unfiltered 10-s frames using a Canon digital camera (+ 200-mm f/3.2 lens) on Feb. 1.862 UT, its position given as α = 16^{h}52^{m}18^{s}.87, δ = −37^{o}54′18′′.9 (equinox J2000.0). The variable was designated PNV J16521887–3754189 when it was posted at the Central Bureau’s TOCP webpage. Additional CCD magnitudes for the variable: Jan. 30.868, 12.5 (Nishimura); Feb. 2.857, 12.1 (Nishimura); 6.836, 12.7 (Nishimura); 10.702, B = 14.7, V = 14.1, R_{c} = 11.8, I_{c} = 12.5 (S. Kiyota, Kamagaya, Japan; remotely with a 0.50-m astrograph at Siding Spring; position end figures 18^{s}.63, 16′′.4); 10.835, 12.6 (T. Noguchi, Chiba-ken, Japan, 0.23-m reflector; position end figures 18^{s}.66, 16′′.7; communicated by Nakano); 23.718, R = 11.7 (S. Foglia and colleagues, using a 0.15-m refractor at Siding Spring; position end figures 18^{s}.65, 16′′.3). Noguchi reports that a star of mag 15.9 is present at the position of PNV J16521887–3754189 on a Digitized Sky Survey image from 1995. P. Schmeer, Bischmisheim, Germany, writes that the red star noted by Noguchi has blue magnitude 18.5 and red magnitude 15.0-15.9 in the USNO-B1.0 catalogue, and V = 17.0 in the GSC2.3 catalogue. Further details are given on CBET 4364.

Optical spectroscopy (range 480-880 nm; resolution 0.56 nm) obtained on Feb. 12.3 UT by J. Strader, L. Chomiuk, A. Bahramian, and C. Britt, Michigan State University, with the 4.1-m SOAR telescope (+ Goodman Spectrograph) at Cerro Pachon shows a spectrum for PNV J16521887– 3754189 that indicates a classification as an “He/N”-type nova. Details are given on CBET 4364.

N. Samus and E. Kazarovets note that this variable has been given the permanent GCVS designation V1657 Sco.

COMET 2P/ENCKE Numerous observers have noticed prominent plasma and dust tails in

recent weeks in images of comet 2P. Images taken by M. Jäger (Vienna, Austria) suggest that the anti-tail has been especially prominent in the past month, fading somewhat since late January. Visual total-magnitude and coma-diameter estimates: 2016 Dec. 26.82 UT, 11.5, 4′ (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 31.72, 13.8, 2′.0 (W. Hasubick, Buchloe, Germany, 44-cm reflector); 2017 Jan. 19.79, 11.1, 4′ (Gonzalez); 24.39, 12.2, 1′.7 (S. Yoshida, Ibaraki, Japan, 0.40-m reflector); 28.75, 12.8, 1′.1 (Hasubick); Feb. 16.77, 9.2, 2′.1 (Hasubick); 16.82, 9.2, 3′ (Gonzalez, 25×100 binoculars); 18.43, 9.0, 4′ (K. Yoshimoto, Yamaguchi, Japan, 26×100 binoculars).

2017 February 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2017 CBAT  Daniel W. E. Green